OpenStudy (anonymous):

write an equation in slope intercept form,y+x=3 then graph it

OpenStudy (gorv):

y=-x+3 slope = -1 y intercept=3

OpenStudy (gorv):

slope intercept form y=mx+c m=slope c= intercept @BOOKER23

OpenStudy (anonymous):

bro its y+x=3 and the formula is y=mx=b

OpenStudy (gorv):

no its y=mx+b

OpenStudy (gorv):

in place of b u can take any other than x,y, and m as they are already constant

OpenStudy (anonymous):

thanks bruh